OpenStudy (anonymous):

Evaluate the triple integral SSSzdV where the region is bounded by the planes y=0, z=0,x+y=1 and the cylinder y^2+z^2=1 in the first octant

OpenStudy (cruffo):

so we note that 0 <=y <=1-x and 0 <=z <=sqrt(1-y^2) So we can deduce that 0 < = x <=1 So far so good???

OpenStudy (cruffo):

ok...continuing \[\int_0^1 \int_0^{1-x} \int_0^{\sqrt{1-y^2}} z \;dz \;dy\;dx\]

OpenStudy (anonymous):

\[1/2 \int\limits\limits_{0}^{1}\int\limits_{0}^{1-x}1-y ^{2} dydx\]

OpenStudy (cruffo):

good...

OpenStudy (anonymous):

\[1/2\int\limits_{0}^{1}(1-x)-((1-x)^{3}/3)dx ?\]

OpenStudy (cruffo):

good...

OpenStudy (cruffo):

The first two terms are easier to deal with, \[\frac{1}{2} \int (1 - x) \;dx = \frac{1}{2}x-\frac{1}{4}x^2\] and you can use u-substitution on (1-x)^3 u = 1-x so du = -dx \[\frac{1}{6} \int (1 - x)^3 \;dx = -\frac{1}{6} \int u^3 \;du = -\frac{1}{24} (1 - x)^4 \] So all together, \[\frac{1}{2}x-\frac{1}{4}x^2 - (-\frac{1}{24} (1 - x)^4)\] \[ = \frac{1}{2}x-\frac{1}{4}x^2 +\frac{1}{24} (1 - x)^4\] Then plug in the limits of integration.

OpenStudy (anonymous):

final answer is 5/24?

OpenStudy (cruffo):

1/4 - 1/24 = 5/24 .... yep that's what I got. did you draw the region? looks like a weird wedge of cheese :)

OpenStudy (anonymous):

haha yes it does, thanks dude i sorta just needed someone to make sure the same answer cheers!